# Representation Shattering in Transformers: A Synthetic Study with Knowledge Editing

### v3

@@ Abstract
Knowledge Editing (KE) algorithms alter models’ weights to perform targeted updates to incorrect, outdated, or otherwise unwanted factual associations. To better identify the possibilities and limitations of these approaches, recent work has shown that applying KE can adversely affect models’ factual recall accuracy and diminish their general reasoning abilities. While these studies give broad insights into the potential harms of KE algorithms, e.g., via performance evaluations on benchmarks, we argue little is understood as to why such destructive failures occur. Is it possible KE methods distort representations of concepts beyond the targeted fact, hence hampering abilities at broad? If so, what is the extent of this distortion? Motivated by such questions, we define a novel synthetic task wherein a Transformer is trained from scratch to internalize a “structured” knowledge graph. The structure enforces relationships between entities of the graph, such that editing a factual association has “trickling effects” on other entities in the graph (e.g., altering X’s parent is Y to Z affects who X’s siblings’ parent is). Through evaluations of edited models and analysis of extracted representations, we show that KE inadvertently affects representations of entities beyond the targeted one, distorting relevant structures that allow a model to infer unseen knowledge about an entity. We call this phenomenon representation shattering and demonstrate that it results in degradation of factual recall and reasoning performance more broadly. To corroborate our findings in a more naturalistic setup, we perform preliminary experiments with pre-trained Llama and Mamba models, reproducing the representation shattering effect therein as well. Overall, our work yields a precise mechanistic hypothesis to explain why KE has adverse effects on model abilities.
@@ 1 Introduction
Large language models (LLMs) have led to unprecedented advances in several domains (Gemini Team, 2023; Bubeck et al., 2023; Touvron et al., 2023; Thoppilan et al., 2022; Chowdhery et al., 2022; Qin et al., 2023; Chen et al., 2021; Ahn et al., 2022; Driess et al., 2023). However, the static nature of their training pipelines implies that as our world evolves, models’ internalized knowledge can become incorrect or outdated. To address this, recent work has proposed several protocols for knowledge editing (KE), wherein the goal is to minimally and precisely alter model weights such that only the targeted information (and its relevant associations) are updated, but all unrelated information remains (ideally) unaffected (Mitchell et al., 2022; Meng et al., 2022a; 2023; Dai et al., 2021; Cheng et al., 2023; De Cao et al., 2021; Sinitsin et al., 2020).
Despite significant work on the topic, it still remains unclear precisely what effects KE should have on a model. For example, assume you edit the fact that “Michael Jordan won the 1998 NBA MVP” to “Reggie Miller won the 1998 NBA MVP”, then what should the impact of such an edit be? Should the model now believe Michael Jordan and the Chicago Bulls never reached the NBA finals in 1998? Should it perhaps believe Reggie Miller was on the Chicago Bulls? Should the pop quote “Be like Mike” (Wikipedia, 2024) now become “Be like Reggie”? As Hofweber et al. (2024); Hase et al. (2024) argue, it is difficult to design clear, well-defined answers for such questions. Motivated by this, recent work has started investigating precisely what effects KE actually has on the model (Hoelscher-Obermaier et al., 2023; Li et al., 2023b; Lynch et al., 2024). For example, Cohen et al. (2023) demonstrate that knowledge beyond the edited fact can often be impacted in a detrimental manner, such that the model begins to have an incoherent understanding of the world; Gupta et al. (2024a) demonstrate unrelated facts are often forgotten by the model post-editing; and Gu et al. (2024) show that KE can harm broader reasoning capabilities beyond mere factual recall. While these works clearly demonstrate the detrimental impacts of editing on a model, they still leave open the question precisely why such harms occur—at a mechanistic level, how are model representations impacted such that a broad set of knowledge and capabilities in a model are heavily distorted once an edit occurs?
This work. To address the questions above, we aim to develop a mechanistic understanding of the impact of KE on a model’s internals. For this purpose, we argue we must solve two problems: (i) identify how a model expresses knowledge about some predefined set of entities in its representations, and (ii) investigate how this mechanism is affected as we apply KE to alter a fact corresponding to a subset of the entities. Instead of attacking a complicated system that may be difficult to interpret (e.g., an off-the-shelf LLM), we take inspiration from a multitude of recent papers that establish synthetic abstractions of the target system and develop precise hypotheses as to why the phenomenon-in-question occurs (Allen-Zhu & Li, 2023c; a; b; Okawa et al., 2023; Chan et al., 2022; Li et al., 2023a; Lubana et al., 2024). Specifically, we define a data-generating process that yields entities arranged in a structured knowledge graph. This structure is defined via use of a predefined set of relations that locally constrain how entities relate to each other (similar to parent-child relations). Given enough entities and relations, such local constraints manifest a broader global structure in the knowledge graph. Performing traversal over the nodes of this knowledge graph, we get sequences that can be used as “strings” to train a Transformer on. As we show, this protocol leads to the model precisely encoding the structure of the graph in its latent representations. However, when KE is applied to edit either incorrectly learned facts or insert counterfactual knowledge (using the method proposed by Meng et al. (2022a)), we find latent representations are heavily distorted and the graph structure completely destroyed—we call this phenomenon representation shattering. Interestingly, this phenomenon manifests in proportion to how far the proposed edit moves a given node from its current location to a new location in the graph (defined via edge distance). We thus hypothesize representation shattering underlies the detrimental effects of KE on a pretrained model’s factual and reasoning capabilities at broad. Overall, we make the following contributions in this work.
@@
Figure 1: Representation shattering explains why knowledge editing can sometimes degrade models’ general capabilities. (a) Prior works finds that editing facts, e.g., the president of the US, can sometimes harm general abilities of LLMs (figure reproduced from Gu et al. (2024)). (b) We introduce a synthetic data generation process defined by a knowledge graph containing ring-shaped geometries. When we train a model on our synthetic data, we observe that the model’s internal representations mirror the ring structure of the underlying data generation process. We then explain the post-edit degradation of the model’s capabilities by uncovering the “shattering” of latent representations. In the example provided, while the edit successfully changes the fact (Entity #124 is to the right of Entity #123), the model’s performance metrics drop after the knowledge edit.
@@
  • Investigations with Off-the-Shelf LLMs. Using pre-trained Llama and Mamba models, we provide evidence for our claims about representation shattering in more naturalistic settings. For one, we find real-world analogues to our synthetic knowledge graph structures (i.e., days of the week) yield similar shattering phenomena in Llama and Mamba models to what we observe in our toy setup (see Sec. 4.5). Additionally, we further reinforce the generality of our findings with preliminary replications of representation shattering under more complex knowledge graph geometries, such as trees (i.e. countries and their cities, Appx. G.3).
@@ 2 Related Work
Knowledge Editing. Several protocols for knowledge editing (KE) have been proposed in recent work. Early work defined meta-learning based approaches (Sinitsin et al., 2020; De Cao et al., 2021; Mitchell et al., 2022) and established the broader desiderata for what properties a KE protocol should satisfy; e.g., ensuring facts unrelated to the target one are not hampered via the editing protocol. Building on work aimed at understanding how Transformers encode knowledge in their internals (Geva et al., 2020), modern KE protocols focus on defining closed-form operations that involve (i) localizing knowledge to specific components in a model (e.g., MLP layers) and (ii) defining operations to alter a factual association by assuming the fact is represented in a localized manner (Meng et al., 2022a; 2023).
@@ 3.1 Knowledge Graphs
Formally, a knowledge graph $G=(X,R,F)$ consists of nodes $X$, relations $R$, and facts $F$, where each fact $f=(x_{i},r,x_{j})\in R$ is defined by a relation $r\in R$ between two entities $x_{i},x_{j}\in X$.
@@
Related pieces of information such as “Alice’s advisor was Bob” and “Bob’s advisor was Carol” can be composed to form cohesive statements such as “Alice’s advisor’s advisor was Carol. To capture such statements, we define compositions of relations below. The composition of relations are essential to capture ripple effects that occur in the knowledge graph after an edit (Cohen et al., 2023) to a relation in $R$.
@@
3.2 Cyclic graphs: a description of the entities and relations
Figure 2: Isomap projections of representations in Llama-3.1-405B (Fiotto-Kaufman et al., 2024). The geometry of the data—for example, the cyclic nature of months or days—is often reflected in the representations learnt by language models. Similar representations can also be found in other models like GPT-2-Small and Mistral 7B (Engels et al., 2024).
We mainly study knowledge graphs where every relation sub-graph is a set of disjoint cyclic graphs, i.e., for every entity $x_{i}$ and relation $r$, there exists exactly one entity $x_{j}$ such that $(x_{i},r,x_{j})\in F$. We specifically choose a cyclic geometry as a global constraint on the graph structure since cycles are a common pattern that relate entities in natural language domains; e.g., see Fig. 2, where we show a 2D projection of representations from Llama-3.1-405B corresponding to months of a year and days of the week naturally organize in a cyclic fashion. For supplementary explorations of geometries of concepts other than cycles (i.e. trees), refer to Appx. G.3.
@@ 3.2 Cyclic graphs: a description of the entities and relations
Depending on the fact being edited, the 3 cyclic orders are used to define the edit sub-graph, the retain sub-graph, and the test sub-graph. We need 3 cyclic orders because knowledge editing methods target edit sub-graph relations; the facts based on these edit relations are then tested to check if a knowledge edit was successful. Meanwhile, the retain sub-graph relations are used by the knowledge editing algorithm to minimize changes to unrelated relations, but no edits are made to facts that use these relations. Finally, the test sub-graph relations are used to define facts that are neither directly edited, nor used by the knowledge editing algorithm. The relations are used to evaluate whether unrelated facts remain unchanged after a knowledge edit. We note that relations for all 3 sub-graphs are seen during pre-training and this distinction between the cyclic orders is made only during model editing. Finally, we note the distance of an edit (shown in Fig. 3) is defined as the shortest distance between the original and edited entity in the cyclic order.
@@ 3.3 Experimental Setup
Data-generating process. We generate a sequence of alternating entities and relations resembling $x_{1}\vec{r}_{1}x_{2}\vec{r}_{2}x_{3}\vec{r}_{3}\dots$, where any consecutive triplet of entity, relation, and entity $x_{i}r_{i}x_{i+1}$ from the sequence is a fact $(x_{i},\vec{r}_{i},x_{i+1})$ in the knowledge graph. The composition of relations $\vec{r}_{i}=r_{i1}r_{i2}r_{i3}\dots$ is a sequence of 1 or more relation tokens, while $x_{i}$ is a single entity token. Every token is sampled using a uniform probability over all the permissible choices (see Alg. 1). For example, a plausible sequence for the example in Fig. 3 is “1 I_C4 4 III_A2 8 III_A3 3 II_C2 7”, which is an alternating sequence of entities and k-hop relations. As previously noted, relations belonging to all three cyclic orders are included in the data generation process; the distinction between edit, retain, and test relations is only relevant to knowledge editing on a trained model. Furthermore, we remark that this sampling process is identical to traversing random walks on the knowledge graph, similar to previous works (Prystawski et al., 2024; Khona et al., 2024). Additional details of the generation process are documented in Appx. B.
@@
Training setup. We train a Transformer model using next-token prediction on the synthetic data generated from the above data generation process. For all experiments (unless stated otherwise), we use a 2-layer nanoGPT Transformer (Karpathy, 2021). For additional details about the toy model, see Appx. C. For experiments with real-world LLMs, see Sec. 4.5.
@@
Evaluation (seen facts). We assess the model’s ability to remember facts seen during training, both before and after an edit. Specifically, to analyze whether the model has learned the fact $(x_{i},\vec{r},x_{j})$, we prompt it with an entity $x_{i}$ and a relation $\vec{r}$, expecting it to produce $x_{j}$ as the next token. In practice, the model outputs can vary across prompts: we account for this by averaging the softmax probabilities across $5$ randomly sampled sequences of the form ctx $x_{i}\vec{r}$ and using the output token with the highest probability. Herein, ctx $x_{i}$ denotes a randomly sampled sequence that ends in $x_{i}$.
@@
Evaluation (unseen facts). We also evaluate the model on two other criteria. (1) Compositional inference. In addition to facts seen in the training data, we evaluate the model on compositions of relations. The model must preserve geometric structures of the data in order to compositionally generalize after a knowledge edit. (2) Logical inference. A key feature of reasoning in natural language is logical inference. For example, if Alice is said to be the advisor of Bob, then Bob is an advisee of Alice (even if it is not explicitly stated). Our data generation process has similar relations, such as clockwise and anti-clockwise 1-hop neighbors. By “holding out” one direction for some such pairs of relations from being observed verbatim in the training dataset, i.e., the relation may only appear compositionally, we can assess the degree to which the model internalizes properties among related relations. We can also evaluate if editing a fact for a relation changes the fact for other related relations, i.e., we check if the model’s knowledge is logically self-consistent after an edit.

### v5

@@ Abstract
Knowledge Editing (KE) algorithms alter models’ weights to perform targeted updates to incorrect, outdated, or otherwise unwanted factual associations. However, recent work has shown that applying KE can adversely affect models’ broader factual recall accuracy and diminish their reasoning abilities. Although these studies give insights into the potential harms of KE algorithms, e.g., performance evaluations on benchmarks, little is understood about why such destructive failures occur. Motivated by this, we define a novel synthetic task in which a Transformer is trained from scratch to internalize a “structured” knowledge graph. The structure enforces relationships between entities of the graph, such that editing a factual association has “trickling effects” on other entities (e.g., altering X’s parent is Y to Z affects who X’s siblings’ parent is). Through evaluations of edited models on this task, we show that KE inadvertently affects representations of entities beyond the targeted one, distorting relevant structures that allow a model to infer unseen knowledge about an entity. We call this phenomenon representation shattering and demonstrate that it degrades models’ factual recall and reasoning performance. We further corroborate our findings in naturalistic settings with pre-trained Llama and Mamba models as well. Overall, our work yields a precise mechanistic hypothesis to explain why KE has adverse effects on model abilities.
Knowledge Editing, Representation Learning, Transformers, Neural Networks
[…]
@@ 1 Introduction
Figure 1: Representation shattering as a mechanistic hypothesis for why knowledge editing has adverse effects on models’ general capabilities. (a) Prior works find that editing facts, e.g., the president of the US, can harm general abilities of LLMs (figure reproduced from Gu et al. (2024)). (b) We introduce a synthetic data generation process (DGP) defined by a knowledge graph containing ring-shaped geometries. Training a model on this data, we observe that the model’s internal representations mirror the ring structure of the underlying DGP. We explain the post-edit degradation of the model’s capabilities by uncovering the “shattering” of latent representations. E.g., in the provided illustration, while the edit successfully changes the fact (Entity #124 is to the right of Entity #123), it also degrades model’s broader performance.
Large language models (LLMs) have led to unprecedented advances in several domains (Gemini Team, 2023; Bubeck et al., 2023; Touvron et al., 2023; Thoppilan et al., 2022; Chowdhery et al., 2022; Qin et al., 2023; Chen et al., 2021; Ahn et al., 2022; Driess et al., 2023). However, the static nature of their training pipelines implies that as our world evolves, models’ internalized knowledge can become incorrect or outdated. To address this, recent work has proposed several protocols for knowledge editing (KE), wherein the goal is to minimally and precisely alter model weights such that only the targeted information (and its relevant associations) are updated, but all unrelated information remains (ideally) unaffected (Mitchell et al., 2022; Meng et al., 2022a, 2023; Dai et al., 2021; Cheng et al., 2023; De Cao et al., 2021; Sinitsin et al., 2020).
Despite significant work on the topic, it still remains unclear precisely what effects KE should have on a model. For example, assume you edit the fact that “Michael Jordan won the 1998 NBA MVP” to “Reggie Miller won the 1998 NBA MVP”, then what should the impact of such an edit be? Should the model now believe Michael Jordan and the Chicago Bulls never reached the NBA finals in 1998? Should it perhaps believe Reggie Miller was on the Chicago Bulls? Should the pop quote “Be like Mike” (Wikipedia, 2024) now become “Be like Reggie”? As Hofweber et al. (2024); Hase et al. (2024) argue, it is difficult to design clear, well-defined answers for such questions. Motivated by this, recent work has started investigating precisely what effects KE actually has on the model (Hoelscher-Obermaier et al., 2023; Li et al., 2023b; Lynch et al., 2024). For example, Cohen et al. (2023) demonstrate that knowledge beyond the edited fact can often be impacted such that the model begins to have an incoherent understanding of the world; Gupta et al. (2024a) demonstrate unrelated facts are often forgotten by the model post-editing; and Gu et al. (2024) show that KE can harm broader reasoning capabilities beyond mere factual recall. While these works clearly demonstrate the detrimental impacts of editing on a model, they still leave open the question precisely why such harms occur—at a mechanistic level, how are model representations impacted such that a broad set of knowledge and capabilities in a model are heavily distorted once an edit occurs?
This work. To address the questions above, we aim to develop a mechanistic understanding of the impact of KE on a model’s internals. For this purpose, we argue we must solve two problems: (i) identify how a model expresses knowledge about some predefined set of entities in its representations, and (ii) investigate how this mechanism is affected as we apply KE to alter a fact corresponding to a subset of the entities. Instead of attacking a complicated system that may be difficult to interpret (e.g., an off-the-shelf LLM), we take inspiration from a multitude of recent papers that establish synthetic abstractions of the target system and develop precise hypotheses as to why the phenomenon-in-question occurs (Allen-Zhu & Li, 2023c, a, b; Okawa et al., 2023; Chan et al., 2022; Li et al., 2023a; Lubana et al., 2024). Specifically, we define a data-generating process that yields entities arranged in a structured knowledge graph. This structure is defined via use of a predefined set of relations that locally constrain how entities relate to each other (similar to parent-child relations). Given enough entities and relations, such local constraints manifest a broader global structure in the knowledge graph. Performing traversal over the nodes of this knowledge graph, we get sequences that can be used as “strings” to train a Transformer on. As we show, this protocol leads to the model precisely encoding the structure of the graph in its latent representations. However, when KE is applied to edit either incorrectly learned facts or insert counterfactual knowledge (using the method proposed by Meng et al. (2022a)), we find latent representations are heavily distorted and the graph structure completely destroyed—we call this phenomenon representation shattering. Interestingly, this phenomenon manifests in proportion to how far the proposed edit moves a given node from its current location to a new location in the graph (defined via edge distance). We thus hypothesize representation shattering underlies the detrimental effects of KE on a pretrained model’s factual and reasoning capabilities at broad. Overall, we make the following contributions in this work.
@@
  • Investigations with Off-the-Shelf LLMs. Using pre-trained Llama and Mamba models, we provide evidence for our claims about representation shattering in more naturalistic settings. For one, we find real-world analogues to our synthetic knowledge graph structures (i.e., days of the week) yield similar shattering phenomena in Llama and Mamba models to what we observe in our toy setup (see Sec. 4.5). Additionally, we further reinforce the generality of our findings with preliminary replications of representation shattering under more complex knowledge graph geometries, such as trees (i.e. countries and their cities, Appx. H.3).
@@ 2 Related Work
Knowledge Editing. Several protocols for knowledge editing (KE) have been proposed in recent work. Early work defined meta-learning based approaches (Sinitsin et al., 2020; De Cao et al., 2021; Mitchell et al., 2022) and established the broader desiderata for what properties a KE protocol should satisfy; e.g., ensuring facts unrelated to the target one are not hampered via the editing protocol. Building on work aimed at understanding how Transformers encode knowledge in their internals (Geva et al., 2020), modern KE protocols focus on defining closed-form operations that involve (i) localizing knowledge to specific components in a model (e.g., MLP layers) and (ii) defining operations to alter a factual association by assuming the fact is represented in a localized manner (Meng et al., 2022a, 2023).
@@ 3.1 Knowledge Graphs
Formally, a knowledge graph $G=(X,R,F)$ consists of nodes $X$, relations $R$, and facts $F$, where each fact $f=(x_{i},r,x_{j})\in F$ is defined by a relation $r\in R$ between two entities $x_{i},x_{j}\in X$.
@@
Related pieces of information such as “Alice’s advisor was Bob” and “Bob’s advisor was Carol” can be composed to form cohesive statements such as “Alice’s advisor’s advisor was Carol. To capture such statements, we define compositions of relations. The composition of relations are essential to capture ripple effects that occur in the knowledge graph after an edit (Cohen et al., 2023) to a relation in $R$.
@@
3.2 Cyclic Graphs: Description of Entities/Relations
Figure 2: Isomap projections of representations in Llama-3.1-405B (Fiotto-Kaufman et al., 2024). The geometry of the data—for example, the cyclic nature of months or days—is often reflected in the representations learned by language models. Similar representations can also be found in other models like GPT-2-Small and Mistral 7B (Engels et al., 2024).
We mainly study knowledge graphs where every relation sub-graph is a set of disjoint cyclic graphs, i.e., for every entity $x_{i}$ and relation $r$, there exists exactly one entity $x_{j}$ such that $(x_{i},r,x_{j})\in F$. We specifically choose a cyclic geometry as a global constraint on the graph structure since cycles are a common pattern that relate entities in natural language domains; e.g., see Fig. 2, where we show a 2D projection of representations from Llama-3.1-405B corresponding to months of a year and days of the week naturally organize in a cyclic fashion. For supplementary explorations of geometries of concepts other than cycles (i.e. trees), refer to Appx. H.3.
@@ 3.2 Cyclic Graphs: Description of Entities/Relations
Depending on the fact being edited, the 3 cyclic orders are used to define the edit sub-graph, the retain sub-graph, and the test sub-graph. We need 3 cyclic orders because knowledge editing methods target edit sub-graph relations; the facts based on these edit relations are then tested to check if a knowledge edit was successful. Meanwhile, the retain sub-graph relations are used by the knowledge editing algorithm to minimize changes to unrelated relations, but no edits are made to facts that use these relations. Finally, the test sub-graph relations are used to define facts that are neither directly edited, nor used by the knowledge editing algorithm. The relations are used to evaluate whether unrelated facts remain unchanged after a knowledge edit. We note that relations for all 3 sub-graphs are seen during pre-training and this distinction between the cyclic orders is made only during model editing. Finally, the distance of an edit (shown in Fig. 3) is defined as the shortest distance between the original and edited entity in the cyclic order.
@@ 3.3 Experimental Setup
Data generation process. We generate a sequence of alternating entities and relations resembling $x_{1}\vec{r}_{1}x_{2}\vec{r}_{2}x_{3}\vec{r}_{3}\dots$, where any consecutive triplet of entity, relation, and entity $x_{i}r_{i}x_{i+1}$ from the sequence is a fact $(x_{i},\vec{r}_{i},x_{i+1})$ in the knowledge graph. The composition of relations $\vec{r}_{i}=r_{i1}r_{i2}r_{i3}\dots$ is a sequence of 1 or more relation tokens, while $x_{i}$ is a single entity token. Every token is sampled using a uniform probability over all the permissible choices (see Alg. 1). For example, a plausible sequence for the example in Fig. 3 is “1 I_C4 4 III_A2 8 III_A3 3 II_C2 7”, which is an alternating sequence of entities and k-hop relations. As previously noted, relations belonging to all three cyclic orders are included in the data generation process; the distinction between edit, retain, and test relations is only relevant to knowledge editing on a trained model. Furthermore, we remark that this sampling process is identical to traversing random walks on the knowledge graph, similar to previous works (Prystawski et al., 2024; Khona et al., 2024). Additional details of the generation process are documented in Appx. A and Appx. B.
@@
Training setup. We train a Transformer model using next-token prediction on the synthetic data generated from the above data generation process. For all experiments (unless stated otherwise), we use a 2-layer nanoGPT Transformer (Karpathy, 2021). For additional details about the toy model, see Appx. D. For experiments with real-world LLMs, see Sec. 4.5.
@@
Evaluation (seen facts). We assess the model’s ability to remember facts seen during training, both before and after an edit. Specifically, to check if the model has learned the fact $(x_{i},\vec{r},x_{j})$, we prompt it with an entity $x_{i}$ and a relation $\vec{r}$, expecting it to produce $x_{j}$ as the next token. In practice, the model outputs can vary across prompts: we account for this by averaging the softmax probabilities across $5$ randomly sampled sequences of the form ctx $x_{i}\vec{r}$ and using the output token with the highest probability. Herein, ctx $x_{i}$ denotes a randomly sampled sequence that ends in $x_{i}$.
@@
Evaluation (unseen facts). We also evaluate the model on two other criteria. (1) Compositional inference. In addition to facts seen in the training data, we evaluate the model on compositions of relations. The model must preserve geometric structures of the data in order to compositionally generalize after a knowledge edit. (2) Logical inference. A key feature of reasoning in natural language is logical inference. For example, if Alice is said to be the advisor of Bob, then Bob is an advisee of Alice (even if it is not explicitly stated). Our data generation process has similar relations, such as clockwise and anti-clockwise 1-hop neighbors. By “holding out” one direction for some such pairs of relations from being observed verbatim in the training dataset, i.e., the relation may only appear compositionally, we can assess the degree to which the model internalizes properties among related relations. We can also evaluate if editing a fact for a relation changes the fact for other related relations, i.e., we check if the model’s knowledge is logically self-consistent after an edit. Precise prompt formats for both (1) and (2) are documented in Appx. C.